OpenStudy (shep):

descibe the graph of y=1/2x-10 -3 compared to the graph of y=1/x

OpenStudy (davejavous):

y=1/2x-10 -3 is a little ambiguous. DO you mean y = 0.5x -13 or y= 1/(2x) -13 ?

OpenStudy (3mar):

It will be like that \[f(x)=\frac{ 1 }{ 2x-10 }-3\]

OpenStudy (3mar):

Compared to the parent function \[f(x)=\frac{ 1 }{ x }\]

OpenStudy (3mar):

So that is the shifting roles.

OpenStudy (3mar):

\[y=\frac{ 1 }{ ax+b }+c\] a: responsible for stretching and compression of the graph b: responsible for shifting horizontally (left and right) c: responsible for shifting vertically (up and down) and try this graph.It is cool

OpenStudy (davejavous):

I assume that with the level of difficulty, you should have done transformations of graphs? y=1/x is the same as: \[f(x)=\frac{1}{x}\] The second line is therefore the following transformation: \[f(2x-10)−3\] If you have done transformations of graphs then you should be able to describe each transformation.

OpenStudy (3mar):

https://www.desmos.com/calculator/m7zdmpatsz Hope that helps